# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 2. Preliminary
for every continuous $\mathbb{Z}^{d}$-periodic function $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (namely $u\in C^{0}(\mathbb{T}^{d})$).
@@
    The image $T^{-}(C^{0}(\mathbb{T}^{d}))$ of $T^{-}$ is equicontinuous, as shown in [GT11, Lemma 4.2]. Moreover, it constitutes a family of semiconcave functions with linear modulus and the same constant, as demonstrated in [Zav23, Proposition 2.4.10].
@@
    [Zav23, Proposition 1.3.6] For any $u\in C^{0}(\mathbb{T}^{d})$ and $n\in\mathbb{Z}_{+}$,
    $T^{+n}\circ T^{-n}[u]\leq u\leq T^{-n}\circ T^{+n}[u].$
  3. (c)
    As $S$ is continuous and coercive, for any $x\in\mathbb{R}^{d}$, we can find $y\in\mathbb{R}^{d}$ that attains the infimum in the definition of $T^{-}[u](x)$ and the supremum in the definition of $T^{+}[u](x)$. In other words, given $u\in C^{0}(\mathbb{T}^{d})$, for each $y\in\mathbb{R}^{d}$, there exists $x\in\mathbb{R}^{d}$ such that
    $T^{-}[u](y)=u(x)+S(x,y)$
    and for each $x\in\mathbb{R}^{d}$, there exists $y\in\mathbb{R}^{d}$ such that
    $T^{+}[u](x)=u(y)-S(x,y).$
In the next, we will characterize the regularity of $u,T^{-}[u],T^{+}[u]$ in the following Lemmas 2.5, 2.6 and 2.12 whose proofs are given in the appendix.
@@
Assume that $S$ satisfies Hypothesis 1. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
    If $T^{-}[u](y)=u(x)+S(x,y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}u(x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}T^{-}[u](y);$
  2. (b)
    if $T^{+}[u](x)=u(y)-S(x,y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}T^{+}[u](x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}u(y);$
@@
Assume that $S$ satisfies Hypothesis 1. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
  2. (b)
Theorem 2.7 (Discrete Weak K.A.M.).
There exists a unique constant $\overline{S}$ such that the equation $u=T^{-}[u]-\overline{S}$ admits at least one solution $u\in C^{0}(\mathbb{T}^{d})$.
Such functions $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are called discrete weak K.A.M solutions. The constant $\bar{S}$ is called the effective interaction.
@@
$\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},\ \quad u(y)-u(x)\leq S(x,y)-\bar{S},$
or equivalently, if $u\leq T^{-}[u]-\bar{S}$, or equivalently, if $u\geq T^{+}[u]+\bar{S}$.
@@
$u\leq T^{-}[u]-\bar{S}\leq T^{-2}[u]-2\bar{S}\leq T^{-3}[u]-3\bar{S}\leq\cdots$
This implies that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is non-decreasing. Moreover, we can show that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is uniformly bounded [Zav23, Proposition 1.2.3]. Note that (a) of Remark 2.4 states that the sequence $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ is equicontinuous. Therefore, $\{T^{-n}[u]-n\bar{S}\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $u_{-}\in C^{0}(\mathbb{T}^{d})$, which is a discrete weak K.A.M. solution.
Definition 2.10.
Let $u$ be a sub-action. We say that $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is $u$-calibrated if $u(y)-u(x)=S(x,y)-\bar{S}$. A sequence $\{x_{n}\}_{n\in\mathbb{Z}}$ is called $u$-calibrated if $(x_{n},x_{n+1})$ is $u$-calibrated for all $n\in\mathbb{Z}$. The non-strict set $\mathcal{NS}(u)$ of $u$ collects all $u$-calibrated pairs:
$\mathcal{NS}(u):=\{(x,y)\in\mathbb{R}^{2}\mid u(y)-u(x)=S(x,y)-\bar{S}\}.$
Remark 2.11.
Let $u$ be a sub-action. If $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is $u$-calibrated, then we have
  • $u(y)=T^{-}[u](y)-\bar{S}$;
  • $u(x)=T^{+}[u](x)+\bar{S}$;
  • $T^{-}[u](y)=u(x)+S(x,y)$;
  • $T^{+}[u](x)=u(y)-S(x,y)$.
@@
Assume that $S$ satisfies Hypothesis 1. Let $u$ be a sub-action.
  1. (a)
    If $(x,y)\in\mathcal{NS}(T^{-}[u])$, then both $u$ and $T^{-}[u]$ are differentiable at $x$. Moreover, $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=dT^{-}[u](x)=-\partial_{1}S(x,y)$.
  2. (b)
    If $(x,y)\in\mathcal{NS}(T^{+}[u])$, then both $u$ and $T^{+}[u]$ are differentiable at $y$. Moreover, $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)=dT^{+}[u](y)=\partial_{2}S(x,y)$.
Definition 2.13.
The projected Aubry set is defined as
$\mathcal{A}:=\pi_{1}\left(\bigcap_{u\text{ is a sub-action}}\mathcal{NS}(u)\right).$
This may not be the most common definition, but it can be proven to be equivalent to other definitions (see [GT11, Proposition 10.5]). We list below, without proof, some properties of the projected Aubry set $\mathcal{A}$. Interested readers may refer to Garibaldi and Thieullen’s paper [GT11] and Zavidovique’s note [Zav23].
@@
Assume that $S$ satisfies Hypothesis 1.
@@
  6. (vi)
    $\mathcal{A}^{*}$ is an F-invariant set, i.e., $F(\mathcal{A}^{*})=\mathcal{A}^{*}$ and $F^{-1}(\mathcal{A}^{*})=\mathcal{A}^{*}$.
Theorem 2.15.
Assume that $S$ satisfies Hypothesis 1. If $u$ is a discrete weak K.A.M. solution, then we have
  • (1)
    For any $n\geq 1$,
    $F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right)\subset\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
    So one can define the closed, non-empty and $F$-invariant set
    *$\mathcal{I}^{*}(u):=\bigcap_{n\geq 0}F^{-n}\left(\overline{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)}\right).$
  • (2)
    there exists a discrete weak K.A.M. solution $\widetilde{u}_{-}$ such that $\mathcal{A}^{*}=\mathcal{I}^{*}(\widetilde{u}_{-}).$
  • (3)
    moreover,
    $\mathcal{A}^{*}=\bigcap\mathcal{I}^{*}(u),$

### v2

@@ 2.1. A notion of important invariant set
for every bounded $\mathbb{Z}^{d}$-periodic function (not necessarily continuous) $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ (namely $u\in B(\mathbb{T}^{d})$).
@@
    The image $T^{-}(B(\mathbb{T}^{d}))$ of $T^{-}$ is equicontinuous and has uniformly bounded oscillation, as shown in [GT11, Lemma 4.2].
@@
    We can derive directly from the definition that
    $\forall y\in\mathbb{R}^{d},\quad T^{-n}[u](y)=$ $\inf_{\begin{subarray}{c}(y_{-n},\dots,y_{-1})\in(\mathbb{R}^{d})^{n}\\ y_{0}=y\end{subarray}}\left\{v(y_{-n})+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})\right\},$
    $\forall x\in\mathbb{R}^{d},\quad T^{+n}[u](x)=$ $\sup_{\begin{subarray}{c}(x_{1},\dots,x_{n})\in(\mathbb{R}^{d})^{n}\\ x_{0}=x\end{subarray}}\left\{u(x_{n})-\sum_{i=0}^{n-1}S(x_{i},x_{i+1})\right\}.$
Definition 2.3.
Let $u\in C^{0}(\mathbb{T}^{d})$ and $n\in\mathbb{Z}_{+}$.
  1. (1)
    The optimal backward map is the closed multi-valued map
    $\Sigma_{-}[u]:\mathbb{R}^{d}$ $\rightarrow 2^{\mathbb{R}^{d}}$
    $y$ $\mapsto\mathop{\arg\min}_{x\in\mathbb{R}^{d}}\{u(x)+S(x,y)\};$
  2. (2)
    the optimal forward map is the closed multi-valued map
    $\Sigma_{+}[u]:\mathbb{R}^{d}$ $\rightarrow 2^{\mathbb{R}^{d}}$
    $x$ $\mapsto\mathop{\arg\max}_{y\in\mathbb{R}^{d}}\{u(y)-S(x,y)\};$
  3. (3)
    The $n$-step optimal backward map is the multi-valued map
    $\Sigma_{-}^{n}[u]:\mathbb{R}^{d}$ $\rightarrow 2^{\mathbb{R}^{d}}$
    $y$ $\mapsto\pi_{1}\left(\mathop{\arg\min}_{\begin{subarray}{c}(y_{-n},\dots,y_{-1})\in(\mathbb{R}^{d})^{n}\\ y_{0}=y\end{subarray}}\left\{u(y_{-n})+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})\right\}\right);$
  4. (4)
    the $n$-step optimal forward map is the multi-valued map
    $\Sigma_{+}^{n}[u]:\mathbb{R}^{d}$ $\rightarrow 2^{\mathbb{R}^{d}}$
    $x$ $\mapsto\pi_{n}\left(\mathop{\arg\max}_{\begin{subarray}{c}(x_{1},\dots,x_{n})\in(\mathbb{R}^{d})^{n}\\ x_{0}=x\end{subarray}}\left\{u(x_{n})-\sum_{i=0}^{n-1}S(x_{i},x_{i+1})\right\}\right).$
Remark 2.4.
  1. (a)
    Given $u\in C^{0}(\mathbb{T}^{d})$, as $S$ is continuous and coercive (Hypothesis 1), for any $y\in\mathbb{R}^{d}$, we can find $x\in\mathbb{R}^{d}$ that attains the infimum in the definition of $T^{-}[u](y)$. Similarly, for any $x\in\mathbb{R}^{d}$, we can find $y\in\mathbb{R}^{d}$ that attains the supremum in the definition of $T^{+}[u](x)$. In other words, $\Sigma_{-}[u](y)\neq\emptyset,\forall y\in\mathbb{R}^{d}$ and $\Sigma_{+}[u](x)\neq\emptyset,\forall x\in\mathbb{R}^{d}$.
  2. (b)
    We can derive directly from the definition and Remark 2.2 (b) that for any $u\in C^{0}(\mathbb{T}^{d})$, $m\in\mathbb{Z}_{+}$, $n\in\mathbb{Z}_{+}$ and $y\in\mathbb{R}^{d}$,
    $\Sigma_{-}^{m}[u]\circ\Sigma_{-}^{n}[T^{-m}[u]](y)=\Sigma_{-}^{m+n}[u](y).$
@@
Let $u=T^{-n}[v]$, $v\in C^{0}(\mathbb{T}^{d})$ and $x,y\in\mathbb{R}^{d}$. The following two statements are equivalent:
  1. (i)
    $x\in\Sigma_{-}^{n}[v](y)$;
  2. (ii)
    $y\in\Sigma_{+}^{n}[u](x)$ and $T^{+n}\circ T^{-n}[v](x)=v(x).$
Proof.
(i) $\Rightarrow$ (ii): Suppose that $x\in\Sigma_{-}^{n}[v](y)$. Using the definition of $T^{-}$, by $u=T^{-n}[v]$, we have
(1) $u(y)=\inf_{\begin{subarray}{c}(y_{-n},\dots,y_{-1})\in(\mathbb{R}^{d})^{n}\\ y_{0}=y\end{subarray}}\left\{v(y_{-n})+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})\right\}=v(x)+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})$
where $y_{-n}=x,y_{0}=y$ and for any $(z_{-n+1},\dots,z_{0}=z)\in(\mathbb{R}^{d})^{n}$,
$u(z)=\inf_{\begin{subarray}{c}(z_{-n},\dots,z_{-1})\in(\mathbb{R}^{d})^{n}\\ z_{0}=z\end{subarray}}\left\{v(z_{-n})+\sum_{i={-n}}^{-1}S(z_{i},z_{i+1})\right\}\leq v(x)+\sum_{i={-n}}^{-1}S(z_{i},z_{i+1}),$
where $z_{-n}=x,z_{0}=z$. By combining these equation and inequality and eliminating $v(x)$, we obtain
$u(y)-\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})\geq u(z)-\sum_{i={-n}}^{-1}S(z_{i},z_{i+1}),\quad\forall(z_{-n+1},\dots,z_{0}=z)\in(\mathbb{R}^{d})^{n},$
which implies $T^{+n}[u](x)=u(y)-\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})$ and $y\in\Sigma_{+}^{n}[u](x)$. Notice that the equation (1) says $u(y)-\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})=v(x)$. Hence $T^{+n}[u](x)=v(x)$.
(ii) $\Rightarrow$ (i): Suppose that $y\in\Sigma_{+}^{n}[u](x)$ and $T^{+n}\circ T^{-n}[v](x)=v(x).$ Using the definition of $T^{-}$ and $T^{+}$, by $u=T^{-n}[v]$, we have
$u(y)\leq v(y_{-n})+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})$
for any $(y_{-n},\dots,y_{-1})\in(\mathbb{R}^{d})^{n}$, $y_{0}=y$ and
$T^{+n}[u](x)=\sup_{\begin{subarray}{c}(x_{1},\dots,x_{n})\in(\mathbb{R}^{d})^{n}\\ x_{0}=x\end{subarray}}\left\{u(x_{n})-\sum_{i=0}^{n-1}S(x_{i},x_{i+1})\right\}=u(y)-\sum_{i=0}^{n-1}S(x_{i},x_{i+1})$
where $x_{0}=x,x_{n}=y$. By combining these equation and inequality and eliminating $u(y)$, we obtain
$T^{+n}[u](x)+\sum_{i=0}^{n-1}S(x_{i},x_{i+1})\leq v(y_{-n})+\sum_{i={-n}}^{-1}S(y_{i},y_{i+1})$
for any $(y_{-n},\dots,y_{-1})\in(\mathbb{R}^{d})^{n}$, $y_{0}=x_{n}=y$, $x_{0}=x$. Since $T^{+n}[u](x)=T^{+n}\circ T^{-n}[v](x)=v(x)$, we conclude that $x\in\Sigma_{-}^{n}[v](y)$. ∎
@@
If $u\in T^{-}(C^{0}(\mathbb{T}^{d}))$, then $\Sigma_{+}[u]$ is surjective. In other words,
$\bigcup_{x\in\mathbb{R}^{d}}\Sigma_{+}[u](x)=\mathbb{R}^{d}.$
Proof.
Suppose that $u=T^{-}[v]$ for some $v\in C^{0}(\mathbb{T}^{d})$. By Remark 2.4 (a), for any $y\in\mathbb{R}^{d}$, there exists $x\in\mathbb{R}^{d}$ such that $x\in\Sigma_{-}[v](y)$. It follows from Lemma 2.5 that $y\in\Sigma_{+}[u](x)$. ∎
For any $u\in C^{0}(\mathbb{T}^{d})$ and $x\in\mathbb{R}^{d}$, we have
$T^{+}\circ T^{-}[u](x)=\sup_{y\in\mathbb{R}^{d}}\inf_{z\in\mathbb{R}^{d}}u(z)+S(z,y)-S(x,y)\leq\sup_{y\in\mathbb{R}^{d}}u(x)=u(x).$
Similarly, for any $u\in C^{0}(\mathbb{T}^{d})$ and $n\in\mathbb{Z}_{+}$, we have
(2) $T^{+n}\circ T^{-n}[u]\leq\cdots\leq T^{+}\circ T^{-}[u]\leq u\leq T^{-}\circ T^{+}[u]\leq\cdots\leq T^{-n}\circ T^{+n}[u].$
Let $u\in C^{0}(\mathbb{T}^{d})$, we denote by $\mathcal{O}^{n}[u]$ the closed set
$\mathcal{O}^{n}[u]:=\{x\in\mathbb{R}^{d}\mid T^{+n}\circ T^{-n}[u](x)=u(x)\}.$
It follows from the equation (2) that $\mathcal{O}^{n+1}[u]\subset\mathcal{O}^{n}[u]$ for all $n\in\mathbb{Z}_{+}$. It follows from Remark 2.4 (a) and Lemma 2.5 that
$\Sigma_{-}^{n}[u](\mathbb{R}^{d})=\mathcal{O}^{n}[u],\quad\forall n\in\mathbb{Z}_{+}.$
So $\mathcal{O}^{n}[u]$ is nonempty as the image of the multi-valued map $\Sigma_{-}^{n}[u]$. Hence
$\mathcal{O}^{\infty}[u]:=\bigcap_{n\in\mathbb{Z}_{+}}\mathcal{O}^{n}[u]$
is a nonempty closed subset of $\mathbb{R}^{d}$.
The following proposition reveals that $\mathcal{O}^{\infty}[u]$ can, to some extent, be regarded as a closed invariant set (see Proposition 2.24).
Proposition 2.7.
For any $u\in C^{0}(\mathbb{T}^{d})$, $m\in\mathbb{Z}_{+}$, we have
$\Sigma_{-}^{m}[u](\mathcal{O}^{\infty}[T^{-m}[u]])=\mathcal{O}^{\infty}[u].$
Proof.
By Remark 2.4 (b),
$\Sigma_{-}^{m}[u]\circ\Sigma_{-}^{n}[T^{-m}[u]](\mathbb{R}^{d})=\Sigma_{-}^{m+n}[u](\mathbb{R}^{d}),\quad\forall n\in\mathbb{Z}_{+}.$
This implies that
$\Sigma_{-}^{m}[u](\mathcal{O}^{n}[T^{-m}[u]])=\mathcal{O}^{m+n}[u],\quad\forall n\in\mathbb{Z}_{+}.$
Taking the limit as $n\to+\infty$, we conclude that
$\Sigma_{-}^{m}[u](\mathcal{O}^{\infty}[T^{-m}[u]])\subset\mathcal{O}^{\infty}[u].$
Conversely, for any $x\in\mathcal{O}^{\infty}[u]$, by definition, $x\in\mathcal{O}^{m+n}[u]$ for all $n\in\mathbb{Z}_{+}$. Using Remark 2.4 (b) again, we have $x\in\Sigma_{-}^{m}[u](\mathcal{O}^{n}[T^{-m}[u]])$ for all $n\in\mathbb{Z}_{+}$. Therefore, there exists a sequence $y_{n}\in\mathcal{O}^{n}[T^{-m}[u]]$ such that $x\in\Sigma_{-}^{m}[u](y_{n}).$ By Lemma 2.5, $y_{n}\in\Sigma_{+}^{m}[T^{-m}[u]](x)$, which is a compact subset (depending on $x$). Hence, there exists a convergent subsequence $y_{n_{k}}\rightarrow y\in\Sigma_{+}^{m}[T^{-m}[u]](x)$. Since $x\in\mathcal{O}^{\infty}[u]$, by Lemma 2.5 again, we have $x\in\Sigma_{-}^{m}[u](y)$. Furthermore, since for any $n\in\mathbb{Z}_{+}$, $\mathcal{O}^{n+1}[T^{-m}[u]]\subset\mathcal{O}^{n}[T^{-m}[u]]$ and $\mathcal{O}^{n}[T^{-m}[u]]$ is closed, we conclude $y\in\mathcal{O}^{\infty}[T^{-m}[u]]$. Thus,
$\Sigma_{-}^{m}[u](\mathcal{O}^{\infty}[T^{-m}[u]])\supset\mathcal{O}^{\infty}[u].\qed$
The following theorem is well-known:
Theorem 2.8 (Discrete Weak K.A.M. Theorem).
There exists a unique constant $\overline{S}$ such that the equation $v=T^{-}[v]-\overline{S}$ admits at least one solution $v\in C^{0}(\mathbb{T}^{d})$ and the equation $u=T^{+}[u]+\overline{S}$ admits at least one solution $u\in C^{0}(\mathbb{T}^{d})$.
Such functions $v:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are called negative discrete weak K.A.M solutions and such functions $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ are called positive discrete weak K.A.M solutions. The constant $\overline{S}$ is called the effective interaction.
@@
$\forall(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d},\ \quad u(y)-u(x)\leq S(x,y)-\overline{S},$
or equivalently, if $u\leq T^{-}[u]-\overline{S}$, or equivalently, if $u\geq T^{+}[u]+\overline{S}$.
Definition 2.10.
Let $u$ be a sub-action. We say that $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$ is $u$-calibrated if $u(y)-u(x)=S(x,y)-\overline{S}$. A sequence $\{x_{n}\}_{n\in\mathbb{Z}}$ is called $u$-calibrated if $(x_{n},x_{n+1})$ is $u$-calibrated for all $n\in\mathbb{Z}$. The non-strict set $\mathcal{NS}[u]$ of $u$ collects all $u$-calibrated pairs:
$\mathcal{NS}[u]:=\{(x,y)\in\mathbb{R}^{2}\mid u(y)-u(x)=S(x,y)-\overline{S}\}.$
@@
$u\leq T^{-}[u]-\overline{S}\leq T^{-2}[u]-2\overline{S}\leq T^{-3}[u]-3\overline{S}\leq\cdots$
This implies that the sequence $\{T^{-n}[u]-n\overline{S}\}_{n\in\mathbb{Z}_{+}}$ is non-decreasing. Moreover, we can show that the sequence $\{T^{-n}[u]-n\overline{S}\}_{n\in\mathbb{Z}_{+}}$ is uniformly bounded [Zav23, Proposition 1.2.3]. Note that (a) of Remark 2.2 states that the sequence $\{T^{-n}[u]-n\overline{S}\}_{n\in\mathbb{Z}_{+}}$ is equicontinuous. Therefore, $\{T^{-n}[u]-n\overline{S}\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $u_{-}\in C^{0}(\mathbb{T}^{d})$, which is a negative discrete weak K.A.M. solution.
@@
Let $u$ be a sub-action. Set
$u_{-}:=\lim_{n\to+\infty}T^{-n}[u]-n\overline{S},\quad u_{+}:=\lim_{n\to+\infty}T^{+n}[u]+n\overline{S}$
and
$u_{-+}:=\lim_{n\to+\infty}T^{+n}[u_{-}]+n\overline{S},\quad u_{+-}:=\lim_{n\to+\infty}T^{-n}[u_{+}]-n\overline{S}$
Then $u_{+}=u_{-+}$ and $u_{-}=u_{+-}$.
Proof.
We show the first equality. By Remark 2.11, we have $u\leq u_{-}$ and thus $u_{+}\leq u_{-+}$. Letting $\varepsilon(n):=\sup(u_{-}-T^{-n}[u]+n\overline{S})$ for any $n\in\mathbb{Z}_{+}$, then
$u_{-}-\varepsilon(n)\leq T^{-n}[u]-n\overline{S}.$
Then using the definition of $T^{+}$, we have
$T^{+n}[u_{-}]-\varepsilon(n)\leq T^{+n}\circ T^{-n}[u]-n\overline{S}\leq u-n\overline{S}.$
Then for any $m\in\mathbb{Z}_{+}$, we have
$T^{+(n+m)}[u_{-}]+(n+m)\overline{S}-\varepsilon(n)\leq T^{+m}[u]+m\overline{S}.$
Letting $m\to+\infty$, we have $u_{-+}-\varepsilon(n)\leq u_{+}$ for any $n\in\mathbb{Z}_{+}$. Letting $n\to+\infty$, we have $u_{-+}\leq u_{+}$. ∎
@@
Let $u\in C^{0}(\mathbb{T}^{d})$. Then $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $\widetilde{u}_{+}\in C^{0}(\mathbb{T}^{d})$ and $\{T^{-n}\circ T^{+n}[u]\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $\widetilde{u}_{-}\in C^{0}(\mathbb{T}^{d})$. Furthermore, if $u$ is a sub-action, then
$\widetilde{u}_{+}=u_{+}:=\lim_{n\to+\infty}T^{+n}[u]+n\overline{S}$
is a positive discrete weak K.A.M. solution and
$\widetilde{u}_{-}=u_{-}:=\lim_{n\to+\infty}T^{-n}[u]-n\bar{S}$
is a negative discrete weak K.A.M. solution. We call $(u_{-},u_{+})$ a conjugate pair.
Proof of Proposition 2.13.
Let us provide the proof for the positive discrete weak K.A.M. solution and the the rest can be obtain similarly.
Let $u\in C^{0}(\mathbb{T}^{d})$. It follows from the equation (2) that $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ is non-increasing. It follows from item (a) of Remark 2.2 that $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ is equicontinuous. It follows from item (a) of Remark 2.2 that $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ has uniformly bounded oscillation. Combining this with the fact that $\mathcal{O}^{\infty}[u]\neq\emptyset$, we obtain $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ is uniformly bounded. Therefore, $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ converges uniformly to a function $\widetilde{u}_{+}\in C^{0}(\mathbb{T}^{d})$.
Let $u$ be a sub-action. We show that $\widetilde{u}_{+}=u_{-+}:=\lim_{n\to+\infty}T^{+n}[u_{-}]+n\overline{S}$. Then by Lemma 2.12 we get $\widetilde{u}_{+}=u_{+}$. Letting $\varepsilon(n):=\sup(u_{-}-T^{-n}[u]+n\overline{S})$ for any $n\in\mathbb{Z}_{+}$, then
$u_{-}-\varepsilon(n)\leq T^{-n}[u]-n\overline{S}\leq u_{-}+\varepsilon(n).$
Applying the operator $T^{+n}$ to both sides of the inequality, we have
$T^{+n}[u_{-}]-\varepsilon(n)\leq T^{+n}\circ T^{-n}[u]-n\overline{S}\leq T^{+n}[u_{-}]+\varepsilon(n).$
Then
$T^{+n}[u_{-}]+n\overline{S}-\varepsilon(n)\leq T^{+n}\circ T^{-n}[u]\leq T^{+n}[u_{-}]+n\overline{S}+\varepsilon(n).$
Letting $n\to+\infty$, we obtain $\widetilde{u}_{+}=u_{-+}$. ∎
Lemma 2.14.
Let $u$ be a sub-action. Then
$\{x\mid u_{-}(x)=u_{+}(x)\}=\mathcal{O}^{\infty}[u_{-}]\subset O^{\infty}[u].$
Proof.
Since $\{T^{+n}\circ T^{-n}[u_{-}]\}_{n\in\mathbb{Z}_{+}}$ is non-increasing and converges to $u_{+}$, we obtain $\{x\mid u_{-}(x)=u_{+}(x)\}=\mathcal{O}^{\infty}[u_{-}]$ by the definition of $\mathcal{O}^{\infty}[u_{-}]$. Since $u_{-}\geq u$ and $\{T^{+n}\circ T^{-n}[u]\}_{n\in\mathbb{Z}_{+}}$ is also non-increasing and also converges to $u_{+}$, we obtain $\mathcal{O}^{\infty}[u_{-}]\subset O^{\infty}[u]$. ∎
We can directly deduce from the above lemma:
Proposition 2.15.
$\bigcap_{(u_{-},u_{+})\text{ conjugate pair}}\{x\mid u_{-}(x)=u_{+}(x)\}=\bigcap_{u\text{ weak K.A.M.}}\mathcal{O}^{\infty}[u]=\bigcap_{u\text{ sub-action}}\mathcal{O}^{\infty}[u].$
Moreover, we have
Proposition 2.16.
$\bigcap_{u\text{ sub-action}}\mathcal{O}^{\infty}[u]=\bigcap_{u\text{ sub-action}}\pi_{1}(\mathcal{NS}[u]).$
Proof.
Let $u$ be a sub-action. Then we have
$u(x)\geq T^{+}[u](x)+\overline{S}\geq u_{+}(x),\quad\forall x\in\mathbb{R}^{d}.$
If $x\in\mathcal{O}^{\infty}[u]$, by Proposition 2.13 and the definition of $\mathcal{O}^{\infty}[u]$, it follows that $u_{+}(x)=\widetilde{u}_{+}(x)=u(x)$. Thus
$u(x)=T^{+}[u](x)+\overline{S},$
which implies that $x\in\pi_{1}(NS[u])$.
Let $u$ be a sub-action. If $x\in\pi_{1}(\mathcal{NS}[u])$ for any sub-action $u$, then $x\in\pi_{1}(\mathcal{NS}[T^{+n}[u]])$ for any $n\in\mathbb{Z}_{+}$. Consequently,
$u(x)=T^{+n}[u](x)+n\overline{S}$
for any $n\in\mathbb{Z}_{+}$. This implies $u(x)=u_{+}(x)$. Therefore, by Proposition 2.13, $u(x)=\widetilde{u}_{+}(x)$ for any $n\in\mathbb{Z}_{+}$. That is, $x\in\mathcal{O}^{\infty}[u]$. ∎
The following proposition gives equivalent definitions of the projected Aubry set.
Proposition 2.17.
Denote by $\mathcal{A}$ the projected Aubry set. Then
$\mathcal{A}$ $=\bigcap_{(u_{-},u_{+})\text{ conjugate pair}}\{x\mid u_{-}(x)=u_{+}(x)\}$
$=\bigcap_{u\text{ weak K.A.M.}}\mathcal{O}^{\infty}[u]$
$=\bigcap_{u\text{ sub-action}}\mathcal{O}^{\infty}[u]$
$=\bigcap_{u\text{ sub-action}}\pi_{1}(\mathcal{NS}[u])$
$=\pi_{1}\left(\bigcap_{u\text{ sub-action}}\mathcal{NS}[u]\right).$
Proof.
It follows from [GT11, Lemma 10.3] and [GT11, Proposition 10.5] that
$\bigcap_{u\text{ sub-action}}\pi_{1}(\mathcal{NS}[u])=\pi_{1}\left(\bigcap_{u\text{ sub-action}}\mathcal{NS}[u]\right)=\mathcal{A}.\qed$
We list below, without proof, some fundamental properties of the projected Aubry set $\mathcal{A}$ (see for instance [Ber08, GT11, Zav23]).
Proposition 2.18.
Assume that $S$ satisfies Hypothesis 1 and Hypothesis 2.
@@
2.2. Regularity results in discrete weak K.A.M. theory
In Section 2.2, we operate under the assumption that the function $S$ satisfies both Hypothesis 1 and Hypothesis 2.
The image $T^{-}(B(\mathbb{T}^{d}))$ (resp. $T^{+}(B(\mathbb{T}^{d}))$) constitutes a family of semiconcave (resp. semiconvex) functions with linear modulus and the same constant, as demonstrated in [Zav23, Proposition 2.4.10]. In the next, we will characterize the regularity of $u,T^{-}[u],T^{+}[u]$ in the following Lemmas 2.19 and 2.22 whose proofs are given in the appendix.
Lemma 2.19.
Assume that $S$ satisfies Hypothesis 1 and Hypothesis 2. Let $u\in C^{0}(\mathbb{T}^{d})$.
  1. (a)
    If $x\in\Sigma_{-}[u](y)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}u(x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}T^{-}[u](y);$
  2. (b)
    if $y\in\Sigma_{+}[u](x)$ for some $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}^{d}$, then
    $-\partial_{1}S(x,y)$ $\in\nabla^{-}T^{+}[u](x);$
    $\partial_{2}S(x,y)$ $\in\nabla^{+}u(y);$
Lemma 2.20.
If $v\in SC(\mathbb{T}^{d})$, then $\Sigma_{-}[v]$ is injective. In other words, if $\Sigma_{-}[v](y_{1})\cap\Sigma_{-}[v](y_{2})\neq\emptyset$, then $y_{1}=y_{2}$.